2. Vincent is leasing 100 units of market space in zamboanga city. He computed his fixed costscat P320,000 monthly. He is leasing each space at P14,000 monthly. He spends aroundvP2,950 per rented space monthly for the maintenance. How many spaces must be rented monthly to breakeven? If vincent is targeting a minimum monthly profit of P500,000, how many spaces must be leased?


3. Jane manufactures soft leathers. She spends around P315 per square foot of leather produced and sells it at P300 per square foot. Also, she knows that her fixed costs amount to P1,400,000. What is the minimum number of square feet sold to be able to breakeven? If she only sold 6,000 square feet of leather, did she earn profit? If yes, how much? If no, how much money did she lose?

1
Expert's answer
2021-11-09T01:08:10-0500

2. Fixed cost = P 320000

Leasing revenue = P 14000

Leasing cost = P 2950

Let n = Number of units for breakeven

Step 2

For breakeven,

Total revenue = Total cost

14000n = 2950n + 320000

n = 28.96 ≈

≈ 29 units

Hence, for breakeven, 29 spaces should be rented.

 

Profit = Revenue - Total cost

500000 = 14000n - (2950n + 320000)

500000 = 14000n - 2950n - 320000

n = 74.21 ≈

≈ 74 units 

Hence, for profit of P 500000, 74 spaces should be rented.

3.

She spends 315 per sq, foot of leather

fixed cost 1,400,000

she produced "\frac{1400000}{315}=4444"


now if she sells at 300 then she earns "3000\times4444=1333200"

"loss=1400000-1333200=66800"